y = −4x + 3
16x + 4y = 12
(x,y)=

Answer by sofiyac(983)

y = −4x + 3
16x + 4y = 12
*************
4x+y= 3
16x + 4y = 12 multiply first one by -4 and add to second one
****************
-16x-4y=-12
16x+4y=12 add them
*************
0=0 this is what you get when you add them, which means there's INFINITELY MANY soltuions